Status Report

ISS On-Orbit Status 16 Feb 2002

By SpaceRef Editor
February 16, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

The first of two rest days began with rave reviews of Êyesterday’s onboard EVA Prep Dryrun by the ground. Things had apparently gone so well that “all, EVA community, MCC-H folks and the Capcoms, were thrilled”.

One EVA issue debated by Flight Control was the planned closure of the Node starboard hatch to the Airlock (A/L) during the EVA, which in an emergency would protect for both an alternate ingress path and an incapacitated IV crewmember. In the event of an abort EVA scenario, the Equipment Lock (EL) of the double-chambered A/L would be depressurized. The IV crewmember will then be a closed hatch away from the emergency MPEV (manual pressure equalization valve) in the A/L, but he can probably get there faster than the EV crew. After weighing pros and cons, it was decided to keep the Node hatch closed during the EVA and accept potential risks to some hardware due to vacuum in the EL in the unlikely double-contingency event.

More information on the EVA on Wednesday was uplinked. For example, crew was informed of the desirability of having a video camcorder set up at a DC-1 porthole for viewing the egress of Busch and Walz from the A/L. But this is not a requirement and is left up to crew’s discretion.

The crew performed the regular weekly (Saturday) housecleaning of their station, a 2-hr job for the three residents.

Routine tasks completed today included (a) the daily checkout of the autonomous UF-1 payloads in the Lab (by Dan Bursch), (b) standard maintenance of the Russian SOSH environment control and life support system in the SM (by Carl Walz), and (c) switching the BMP micropurification system’s filter bed #2 Êto regeneration mode, after terminating regeneration of filter #1 earlier in the day (by Yuri Onufrienko).

CDR Onufrienko also again took two control readings, about nine hours apart, of the suspect IK-0501 gas analyzer (GA) in the SM, using the IGZ gas meter. In the last several days, the GA appears to have produced unchanging (static) ppCO2 values at 2.5 mmHg.

The two Flight Engineers used their rest time to complete an item added to the task list last night, namely, to disconnect two umbilicals in the Lab. These DDCU (dc-to-dc converter unit) power output cables had to be demated from two UIPs (utility interface panels) behind closeout panels, one on the Lab’s portside, the other overhead, in support of the DDCU test to be conducted during the EVA (which includes inspection of connections at two DDCUs and their CIDs [circuit interrupt devices] at the Z1 truss).

Tomorrow, 2/17, at 3:21 pm EST, MCC-H plans to maneuver ISS flight attitude from the current LVLH (local vertical local horizontal) reference frame to XPOP (x-axis perpendicular to orbit plane). Since the actual maneuver is performed by Russian segment (RS) thrusters, this requires attitude control handover from the U.S. MCS (motion control system) and its CMGs to the Russian MCS and back.

Science Update (Expedition Four — Ê10th):

Hoffman-Reflex: In progress.

Extra-Vehicular Activity Radiation Monitors (EVARM): The Pre-EVA badge reading was successfully completed this week. Upcoming Pre-EVA badge readings are scheduled for 2/18 and 2/19, and the Post-EVA reading is scheduled for 2/21.

Ultrasound: FE-2 Dan Bursch was thanked for his successful checkout of the HRF (human research facility) Ultrasound hardware. The ground team received the image files and is looking forward to evaluating the on-orbit ultrasonic image quality.

Pulmonary Function in Flight (PuFF): The PuFF team is looking forward to Dan and Carl’s pre-EVA session on Monday, 2/18 (which will also serve as a monthly session) and a session following the EVA.

Renal (Kidney) Stone Experiment: The Renal team thanked the crew for their participation. The team looks forward to continuing the study.

Interactions (NTXN): Continuing on a weekly schedule, nominally on Wednesdays.

Human Research Facility Workstation (HRF WS): The HRF Workstation Functional was successfully completed. The ground team thanked the crew for their efforts and patience with this difficult cabling operation.

Human Research Facility/PC (HRF/PC): This week’s downlink sent 1/15 PuFF files, last week’s EVARM files, and Friday’s Ultrasound images down to the PIs on the ground. The PIs are looking forward to analyzing this data.

Cellular Biotechnology Support Systems (CBOSS): The BTR (biotechnology refrigerator) was keeping samples from the experiment at 4C for the duration of the increment. However, during the five-hour power outage on 2/4 (Monday) there was an excursion to 13C. The impact is not yet known. Based on the data from the manufacturer of the preservative for the samples destined for gene array analysis, specialists anticipate there may be some sample loss. But this too is speculative.

Physics of Colloids in Space (EXPPCS): EXPPCS accomplished four activities during its one 72-hr run this past week. Because the slow fractal sample is aggregating incredibly slowly, the science team elected to do a series of measurements on other samples and return to the examination of the slow fractal at a later time. First, a high resolution measurement of the scattering from the AB13 and AB6 binaries over a range of length scales largely neglected in previous measurements detected a large number of Bragg scattering peaks. Measurements on earth have never revealed so many Bragg peaks on similar samples. Second, routine checkup measurements were performed on the binary colloidal crystals and the colloid polymer crystal samples. These measurements serve to characterize the long term evolution of these samples. Third, a new time-lapse movie of the colloidal glass sample was taken after using a different motion to homogenize the sample. Finally, the ground commanded open the valve separating the colloid and salt solutions on the fast fractal sample. The sample was initiated on 2/15 (Friday).

Space Acceleration Measurement System (SAMS): Currently all five sensors are actively transmitting acceleration data. Will continue to support general characterization in the vibratory regime of the microgravity environment.

Microgravity Acceleration Measurement System (MAMS): MAMS Êcontinues collection and archival of low-frequency acceleration data for the characterization of the ISS quasi-steady microgravity environment. HiRAP has been Êdisabled due to continuing coverage by SAMS and to reduce the downlink bandwidth utilized by acceleration measurements.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): The experiment is running without any problems.

Materials ISS Experiment (MISSE): Nominal and collecting data. The ground is looking forward to viewing the photos of MISSE.

Educational Payload Operations (EPO): EPO will most like be rescheduled for Week 11 (was cancelled on 2/4 because of the RS computer problem and power outage).

Active Rack Isolation System – Isolation Characterization Experiment (ARIS-ICE): The ARIS-ICE team has developed a recovery plan for ARIS. The new crew procedures required are in work. It is hoped to complete the recovery during week 12.

EarthKAM: Then ground team thanked the crew for all their “hard work and support for ISS EarthKAM during our most recent operations! ÊEarthKAM had an incredibly successful mission. To the delight of our middle school students and teachers across the world, we have more than 800 beautiful Earth images taken over the course of our 5-day mission. The students are especially delighted with the pictures taken with the 180mm lensthe detail and clarity observable in these images is truly astounding and beautiful”. The members of the ISS EarthKAM team at UCSD and JSC offered the crew sincere thanks for all their support and dedication that helped make this mission a success. All EarthKAM images are available for public access on the EarthKAM data system at:

Advanced Astroculture (ADVASC): ADVASC was successfully activated on the morning of 2/13 (Wednesday). After some planned activation commanding, the payload settled into the desired set points. The growth chamber footage shows no visual signs of seed germination. ADVASC is in a nominal state.

Zeolite Crystal Growth (ZCG): ZCG team was very excited to find out that the furnace worked as expected during the check-out activity. Furnace heat-up, telemetry reception, and ground commanding were all nominal. Hardware is ready to process samples to be launched on ISS 8A.

CEO (crew earth observations) targets (optional) for today were Angolan Biomass Burning (another good, fair-weather pass for oblique and limb shots, either side of track, to detect early fire season plumes), Nile River Delta (this pass offered nadir views of the city of Cairo and the Suez Canal. Crew was to use the ESC [electronic still camera] to capture the detail and infrastructure of this megacity and vital waterway), European Smog (pass was ideally timed and positioned to view, not smog, but Saharan Dust being swept northeastward in advance of a major winter storm over Western Europe. Of interest: looking to the right of track toward Crete and Greece for a dust plume visible in satellite imagery), Tuamotu Archipelago (there was some cloudiness this pass, but there should have been enough openings to view this high-density portion of the archipelago. Use of ESC to map details of the numerous reefs and atolls there), and Lakes of the Eastern Sierra Watershed (crew was to use this pass to map, document and confirm the reported heavy snow pack from recent snows over the Sierras and the Basin and Range Province further east).

ISS Orbit (as of this morning, 6:20am EST):

  • Mean altitude — 382.9 km
  • Apogee — 386.5 km
  • Perigee — 379.4 km
  • Period — 92.2 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0005276
  • Orbits per 24-hr. day — 15.62
  • Decay rate — 380(mean) in last 24 hours
  • Revolutions since FGB/Zarya launch (Nov. â98) — 18519
  • Current Flight Attitude — LVLH +XVV ZLV (local vertical/local horizontal: +X-axis in velocity vector; Z-axis in local vertical), with TEA (torque equilibrium attitude [pitch: -10 deg, yaw: -7 deg, roll: 0 deg]). Until 2/17.

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.